Show that v(x, y;a, b)= (x+y)[2xy+(a-b)(x-y)+2ab]/(a+b)3 is the Reimann function for second order pde uxy + (2/x+y)(ux + uy)=0


1
Expert's answer
2021-02-02T05:52:24-0500

"L(u)=u_{xy}+a(x,y)u_x+b(x,y)u_y+c(x,y)u=f(x,y)"

"L^*(v)=v_{xy}-(av)_x-(bv)_y+cv"

Condition of Riemann's function:

"L^*(v)=0"


"v_x=\frac{2xy+(a-b)(x-y)+2ab}{(a+b)^3}+\frac{(x+y)[2y+(a-b)]}{(a+b)^3}"

"v_{xy}=\frac{2x-(a-b)}{(a+b)^3}+\frac{2y+(a-b)}{(a+b)^3}+\frac{2(x+y)}{(a+b)^3}=\frac{4(x+y)}{(a+b)^3}"

"a(x,y)v=b(x,y)v=2/(x+y)\cdot (x+y)[2xy+(a-b)(x-y)+2ab]/(a+b)^3="

"=2[2xy+(a-b)(x-y)+2ab]\/(a+b)^3"

"cv=0"

"(av)_x=\frac{2(2y+(a-b))}{(a+b)^3}"

"(bv)_y=\frac{2(2x-(a-b))}{(a+b)^3}"

"L^*(v)=\frac{4(x+y)}{(a+b)^3}-\frac{2(2y+(a-b))}{(a+b)^3}-\frac{2(2x-(a-b))}{(a+b)^3}=0"